OpenStudy (anonymous):

Using the formula x = 2a + y, solve for y Solve the following inequality: 13 – 2x ≥ 7 Solve using the addition principle: -8 + y = 17 y=

OpenStudy (texaschic101):

x = 2a + y you want y by itself, so subtract 2a from both sides x - 2a = y 13 - 2x >= 7 -2x >= 7 - 13 -2x >= - 6 (divide by -2 to get x by itself, but remember, when dividing/multiplying by a negative, changes the sign x <= 3 -8 + y = 17 add 8 to both sides -8 + 8 + y = 17 + 8 (what you do on one side of the equal sign, you have to do on the other side) Now combine like terms y = 17 + 8 y = 25

OpenStudy (texaschic101):

any questions ?

OpenStudy (anonymous):

So 1st one x=2a I always get confused when adding things to each side. it goes on both sides of the = sign correct ?

OpenStudy (anonymous):

Sorry I meant y=2a for the 1st one ..